OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a red marble, do not replace it, and pick a yellow marble?

OpenStudy (anonymous):

choices are 2/15 1/15 1/10

OpenStudy (anonymous):

probability that the first one chosen is red is \(\frac{3}{10}\) probability that the second one is yellow given the first one is red is \(\frac{3}{9}=\frac{1}{3}\) multipliy \[\frac{3}{10}\times \frac{1}{3}\] for your answer

OpenStudy (anonymous):

so 1/10? I have a ton more can you help me with them?

OpenStudy (anonymous):

I think I may be able to figure out everything else thanks! If I don't understnad some I'll put them u!!